OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $8 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Sandlers paid a total of $194 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 194 = one-halfx + 2(x - 8) What is the regular price of a child's ticket? $68 $76 $84 $105

OpenStudy (adajiamcneal):

B

OpenStudy (cj49):

194=1/2(x+ 4(x-8))

OpenStudy (anonymous):

wich one and why

OpenStudy (adajiamcneal):

frac{ 1 }{ 2 }x + 2(x - 8) = 194 Lets distribute that 2 into the parenthesis \frac{ 1 }{ 2 }x + 2x - 16 = 194 Now combine like terms \frac{ 5 }{ 2 }x -16 = 194 Add 16 to both sides \frac{ 5 }{ 2 }x = 210 and finally multiply both sides by 2/5 x = 210 \times \frac{ 2 }{ 5 } so what is 210 times 2? 420 and what is 420 divided by 5? 84 so it would be $84 dollars for an adults ticket....so since a child's ticket is $8 less than that we get 84 - 8 = 76

OpenStudy (anonymous):

thanks cutie hahah

OpenStudy (cj49):

388=x+4x-32 388+32=5x 420=5x x=420/5 x=84 84-8=76

OpenStudy (adajiamcneal):

welcome